Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day08


Crew 99 Log Book for February 5, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: EVA # 12 16:15-16:45

Brief Narrative of Field Mission Results: Tested EVA planner

EVA Data/Interpretations: As per EVA report

Engineering/Hab Maintenance: Nothing major to report.

Report Transmission Schedule: Engineering, EVA #12, journalist's,
chef's, commander's. By 8:30pm

Plans for Tomorrow: Celebrating Superbowl Sunday by not watching the
Steelers/Packers game because we're on Mars.

Inventory: Will inventory remaining dehydrated vegetables this evening

Support Requested: Due to high winds overnight/early this morning, one
of the ATV covers blew away in spite of placing several rocks on it to
hold it down. Any better ideas for keeping those in place?

Miscellaneous: First contact with a group of mysterious "Hanksvillers"
on ATVs. The Martian atmosphere apparently had no ill effects on their
physiology.




Commander's Report

Donna Viola Reporting


Last night, the astronomers worked on realigning the telescope - and

also took some time to teach the rest of the team a little bit about astronomy and the CCD cameras (apparently, it's like a "bucket brigade"). Last night, we also ran into some major problems with our LAMBDA experiment - so today, we ended up cutting off our first trial, reengineering some components, and setting up for a second attempt.


Around lunchtime, we heard a strange sound outside - and peered out the window to see five people on ATVs.


Trying not to be rude, we opted to break sim briefly to greet them. We discovered that they were from this mysterious place called "Hanksville", and had a fascinating discussion about the building of the pyramids and the role of evidence in proving and disproving the existence of highly-evolved reptiles. In the afternoon, Andie taught me how to use the EVA planner, and we went to explore the Thielges Flats.


Max and Sukrit rehydrated dinner, and we are currently wrapping up our celebration of World Nutella Day (if you haven't seen the webcams lately, you may want to check that out before it's too late). Tonight looks like it will be a relaxing night; hopefully, that will appease the crew to the point that they will NOT mutiny within the next week.


Engineering Report

Kevin Newman Reporting


Generator/Electricity:


· Kitty (Diesel Generator) (Used? Yes/No): Yes


· Honey (Gas Generator) (Used? Yes/No): No


· Battery Charge Level:


o 10.59 V everstart


o 12.37 V armor plate


o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is well


Fuel Status:


· Diesel (Full/0.75/0.5/0.25/Empty): 0.25


· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75


· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 20 Gallons


Notes/Comments:


Oil:


· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG


· Honey Oil Quantity (# of Quarts in storage): Leave it to DG


· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench


· Car Oil Quantity (# of Quarts in storage):


Water Status:


· Outside Potable Water Tank Level (inches from bottom): 13


· Trailer Potable Water Tank Level (inches from bottom): FULL


· Hab Potable Water Tank Level (inches from bottom): 9


· Potable Water Meter Reading: 417168


Notes/Comments:


GreenHab:


· Condition of plants in Tank 1: Nothing


· Condition of plants in Tank 2: Ok


· Condition of plants in Tank 3: Good


· Greenhab notes/ comments:


Observatory:


· Telescope Used? (Yes/no) No



Observatory Notes:


Transportation:



Hab Car (NOT RENTAL)


· Used (yes/no): Yes


· Oil Checked (yes/no): No


· Tire Status: Not checked


· Odometer at end of day: 123611.6


· Notes/Comments on Hab Car:


o Took Hab car to Hanksville to fill up gas



ATV 1 (Opportunity):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


ATV 2 (Spirit):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


ATV 3 (Viking I):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


ATV 4 (Viking II):


· Used: No


· Oil Checked: No


· Fuel Consumed: None


· Tire Status: Not Checked


· Comments:


Heating and Ventilation:


· Thermostat setting upstairs (Farenheit): 66


· Actual temperature upstairs (Farenheit): 72


· Thermostat setting downstairs (Farenheit): 65


· Actual temperature downstairs (Farenheit): 67


Computers/Networking Infrastructure:


Notes/comments on internet/computers?: All good

General Engineering

1. Notes/Comments:


Especially windy today, those stones found oddly placed in a semi-circle 300ft north-east of the Hab sure came in handy to hold down the ATV covers.


EVA 12 Report

Andie Gilkey Reporting

Time: 16:30-17:00

Crew Members: Andie Gilkey, Donna Viola

Site Location: UTM: 518500 4251100
Transit Mode: foot

Objectives: Measure how accurately the mission planner predicts astronaut energy consumption, distance traveled, and EVA time on the way to Thielges Flats.

Results: We were able to successfully measure distance traveled and EVA time.

Lessons Learned: Make sure the heart rate monitor is working before you put on the spacesuits.


Astronomer's Report

Sukrit Ranjan Reporting

Yesterday the astronomers attempted to characterize the drift speed of the telescope with the CCD. We did a 2-star alignment, followed by 4 additional calibration stars.

The alignment was far from perfect -- while the telescope was able to locate the Pleiades reasonably well, Polaris was a bust. We hypothesize at least part of the reason to be misalignment of the equitorial head; in particular, the altitude of the telescope is low by ~.5 degree.

We took a 150s exposure of the Pleiades to characterize drift rate and are currently post processing. Qualitatively, drift is very high -- polar alignment must be off.


Cook's Report

Sukrit Ranjan Reporting

1a) Today was

-A non-cooking day.

1b) Was today a special day celebrated at the main meal with special

food or activity?

Absolutely! Today was World Nutella Day: http://www.nutelladay.com/. We performed interpretative dance and human sculpture poses.

2)  List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Gourmet Reserves Chicken Rotelle

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded)  (Example:  2 servings potato soup stored;

2 servings rice burned/inedible, discarded)

- None, we ate it all.

4)  Recipes for anything more complicated than following package

directions exactly, or rehydrating.

Ingredients
1 cup granulated sugar
3/4 cup butter, melted
2 tablespoons water
2 large eggs
1 teaspoon vanilla extract
1 1/3 cups all-purpose flour
1 1/2 cups double chocolate-flavor Nestle Nesquik powder
1/2 teaspoon baking powder
1/4 teaspoon salt
3/4 cup chocolate chips

Directions
1.Preheat oven to 350°.
2.Grease 9 x 13 baking pan.
3.Combine sugar, butter and water in a large bowl; stir well.
4.Stir in eggs and vanilla.
5..Combine flour, Nesquik, baking powder, and salt in medium bowl; stir in sugar mixture.
6Stir in chocolate chips.
7.Spread into prepared pan.
8.Bake for 20 to 25 minutes or until wooden toothpick inserted in center comes out slightly sticky.
9.Cool in pan on wire rack.
10.Cut into bars.

5)  Comments/questions for Kim and Jean, the food study investigators

- Very filling -- very popular. Many people added salt individually.


Journalist Report

Sukrit Ranjan Reporting

Today was World Nutella Day! The team did interpretative dance/sculpture celebrations with lunch (were you watching the webcams? Bet you wish you had been, huh!). Today also saw a considerable reworking of the LAMBDA apparatus.

The astronomers reduced data and worked on determining the telescope drift rate. This evening, we are preparing a great experiment: Nesquick brownies!